# SWBT: Similarity Weighted Behavior Transformer with the Imperfect Demonstration for Robotic Manipulation

### v1

@@ IV-A Simulation Environments and Datasets
IV-B Evaluation Results in Simulation
TABLE I: Comparisons in terms of success rates of the five tasks on ManiSkill2 benchmark. BC and TF-BC are the baselines that can only use expert demonstrations.
@@ IV-B Evaluation Results in Simulation
SWBT-base 83.6 84.0 44.4 26.0 16.0
SWBT 85.6 88.0 50.4 29.6 28.4
TABLE II: Comparisons in terms of success rates of the five tasks on ManiSkill2 benchmark. All baselines are proposed for the offline IL setting with imperfect demonstrations.
Method PickCube StackCube PickYCB Fill Hang
@@
SWBT 85.6 88.0 50.4 29.6 28.4
@@
We comprehensively compared SWBT to many state-of-the-art offline imitation learning algorithms including Behavior Cloning (BC) [12], Transformer-based Behavior Cloning (TF-BC) [36], Decision Transformer (DT) [20], DemoDICE [15], and Discriminator-Weighted Behavioral Cloning (DWBC) [17]. When training the DT, we set the goal as 1 for expert demonstrations and 0 for imperfect demonstrations and hope this signal can help the transformer network distinguish the behavior mode. We trained all methods for 100k gradient steps and evaluated 50 episodes every 5k steps. The final results are calculated as the average success rates of the last 5 checkpoints over 3 random seeds.
In Table I, we compared SWBT with methods BC and TF-BC that can only use expert demonstrations. To verify the effectiveness of the pre-training process, we trained SWBT-base that only leverages the imperfect demonstrations in the pre-training process. Table I shows that SWBT outperformed all baselines consistently on all 5 tasks by a large margin. These results suggest that the high-quality imperfect demonstrations can significantly help improve the final success rate. In addition, the SWBT-base also improved performance on all tasks compared to BC and TF-BC, which shows that the pre-training process can learn expressive features by combining the powerful transformer and imperfect demonstrations.
Figure 4: Comparisons of different quantity and quality of the imperfect demonstrations. The x-axis represents the success rates of the behavior policies that collect the imperfect demonstrations. The y-axis shows the success rates of SWBT. The green and red lines represent the ratios of the expert and imperfect demonstrations are 1:1 and 1:3, respectively.
Table II compared SWBT with methods DT, DemonDICE, and DWBC that simultaneously use both the expert and imperfect demonstration. We can observe that SWBT still surpassed all baseline methods a lot. It is difficult for DemoDICE and DWBC to assign accurate weights for each imperfect high-dimensional transition without a pertaining process. We also found that the training process of DWBC is very unstable without careful hyperparameter tuning. All of the above reasons lead to a decrease in the success rates of these algorithms, even compared to BC. In contrast, SWBT can improve performance by accurately calculating the quality scores and filtering out high-quality demonstrations.
@@ IV-C Ablation Study in Simulation
Impact of the imperfect demonstration quantity and quality. To verify how the imperfect demonstration quantity and quality influence the final performance of our method, we conducted extensive experiments on the StackCube task. We collected five different level imperfect demonstrations using different behavior policies with success rates of 0.00, 0.23, 0.44, 0.60, and 0.86, respectively. For each level, we collected imperfect demonstrations in two ratios, 1:1 and 1:3, where the former 1 represents the 300 expert demonstrations (i.e., 300 and 900 imperfect demonstrations).
TABLE III: Comparisons of different quality score threshold $\beta$ on the StackCube task.
@@
TABLE IV: Comparisons of different similarity functions on the StackCube task.
@@
Impact of quality score threshold $\beta$. We evaluated the impact of various quality score thresholds $\beta$ on the StackCube task. We collected 300 expert trajectories and 900 imperfect trajectories (i.e., 180k transitions) consisting of three groups of 300 trajectories, collected by three level agents with success rates of 0.0, 0.60, and 0.86, respectively.
As shown in Table III, we can observe that as the quality score threshold value $\beta$ increases, the reserved demonstrations become less and less. The highest success rate on the StackCube task is with $\beta$ as 0.90. When $\beta$ is less than 0.90, the performance increases as $\beta$ increases because more and more low-quality demonstrations are discarded. When $\beta$ is larger than 0.90, we argue that the performance drops as $\beta$ increases because more high-quality demonstrations are discarded, degrading the task performance.
[…]
Figure 5: We conducted a digital twin system [48] following the ManiSkill2 for real-world experiments on the PickCube, StackCube, and PickYCB tasks. Experiments show that SWBT can be applied to real-world applications successfully.
Impact of different similarity calculation functions. We evaluated different combinations for the calculation of the similarity distance. We examined $L_{2}$ distance and $Cosine$ similarity distance with the whole segment features denoted as “Seg.” and the last time step feature denoted as “Last” on the StackCube task. As shown in Table IV, “Last+$L_{2}$” achieved the highest performance, while the results of using cosine similarity are relatively lower. We argue that the last time step of the trajectory segment is the most important step for predicting the next action, and the attention mechanism in the transformer has extracted the historical information.
IV-D Real-World Experiments
We applied SWBT to real-world applications by following the setting in ManiSkill2 benchmark and conducting a digital twin system [48]. We trained the policy in the simulation and deployed the model on the real Franka Panda robot. The input of the model includes 2 RGBD images from the simulator and robot proprioceptive states from the real robot. Then the generated actions were applied to both the simulated and real-world robot arms. As shown in Figure 5, we conducted experiments on three rigid body tasks including PickCube, StackCube, and PickYCB. Experiments demonstrate that SWBT can be applied to real-world applications successfully. Please refer to the supplementary video for the manipulation process and more details.
V Conclusion and Limitation
To the best of our knowledge, we are the first to use both expert and imperfect demonstrations simultaneously to train the robot manipulation policy in an offline imitation learning setting. Concretely, we propose a novel framework, named Similarity Weighted Behavior Transformer (SWBT), that calculates accurate quality scores using a pre-trained transformer and then does weighted behavior cloning with the high-quality imperfect demonstrations. The experiments on the ManiSkill2 benchmark and real-world applications demonstrated that SWBT can efficiently and correctly leverage imperfect demonstrations to boost the final performance.
Although SWBT can effectively utilize imperfect demonstrations, it still has a potential limitation. SWBT is extremely data-intensive due to the transformer architecture, and thus it is challenging to apply SWBT to tasks with sparse data, e.g., few-shot learning.

### v2

@@ IV-A Experiment Setup
TABLE I: Comparisons in terms of success rates of the five tasks on ManiSkill2 benchmark. BC and TF-BC are the baselines that only use expert demonstrations. All other baselines used imperfect demonstrations.
@@
TT [43] 22.0 20.0 12.4 14.8 20.4
BeT [44] 82.6 81.2 34.6 10.6 18.6
PACT [48] 82.0 83.2 38.4 14.6 16.0
MTM [49] 66.6 79.4 32.0 23.2 18.4
SMART [50] 65.4 81.2 25.2 8.8 16.2
@@
ISW-BC [26] 79.2 82.4 45.4 26.2 18.8
SSDF-base 83.6 84.0 44.4 26.0 16.0
SSDF 85.6 88.0 50.4 29.6 28.4
Real-world Experiments. We follow Droid [57] to build our real-world robotic arm environment. As shown in Figure 3, we use a Franka Panda 7-DoF robot arm. The input RGB images $o\in\mathbb{R}^{2*270*480*3}$ are from two external fixed-view Zed 2 stereo cameras. We collect the dataset through human demonstration. For each task, we randomly place the objects within a specified area and collect 30 success trajectories as the expert dataset and 70 failed trajectories as the imperfect dataset. We record the RGB images from two camera views and robot proprioceptive states, e.g., joint position $m\in\mathbb{R}^{7}$. Our model predicts the 6D pose, including position $(x,y,z)$ and rotation $(roll,pitch,yaw)$.
IV-B Evaluation Results
TABLE II: Comparisons in terms of success rates of the five tasks on single-arm Franka robot.
Method Place Open Open Stack Arrange Avg.
Tennis Drawer Box Block Flower
BeT [44] 20 30 25 10 0 17
ACT [31] 45 45 65 20 20 39
Diffusion Policy [58] 35 35 60 25 20 35
ISW-BC [26] 40 45 65 20 15 37
SSDF+ACT 50 55 65 30 25 45
@@ IV-B Evaluation Results
Results on Simulation. We comprehensively compared SSDF to many state-of-the-art offline imitation learning algorithms including 4 types. 1) IL methods that only use expert demonstrations: Behavior Cloning (BC) [14], Transformer-based Behavior Cloning (TF-BC) [45], Trajectory Transformer (TT) [43], and Behavior Transformer (BeT) [44]. 2) Pretraining methods that pre-train model with mixed data and then finetune using expert data: Perception-Action Causal Transformer (PACT) [48], Masked Trajectory Models (MTM) [49], and Self-supervised Multi-task pretrAining with contRol Transformer (SMART) [50]. 3) Goal-conditioned methods that take task completion signal as input: Decision Transformer (DT) [46], and Conditional Behavior Transformers (C-BeT) [47]. 4) IL methods that rank imperfect data: DemoDICE [23], Discriminator-Weighted Behavioral Cloning (DWBC) [24], and Importance Sampling Weighted Behavior Cloning (ISW-BC) [26]. To verify the effectiveness of the pre-training process, we trained SSDF-base that only leverages the imperfect demonstrations in the pre-training process. We trained all methods for 100k gradient steps and evaluated 50 episodes every 5k steps. The final results are calculated as the average success rates of the last 5 checkpoints over 3 random seeds.
Table I shows that SSDF outperformed all baselines consistently on all 5 tasks by a large margin. For IL methods that rank imperfect data [23, 24, 26], once they assign high weights to low-quality data points, it in turn interferes with the results of the training. For example, DWBC only achieved a success rate of 32.8 on the PickCube task. This is due to the difficulty of the discriminator in predicting accurate weights for high-dimensional image inputs, which drastically reduced the success rate. In contrast, SSDF can improve performance by accurately calculating the quality scores and filtering out high-quality demonstrations.
Results on Single-arm Franka Robot. For the single-arm franka robotic tasks, we compared 4 state-of-the-art methods that are BeT [44], Action Chunking with Transformers (ACT) [31], Diffusion Policy [58], and ISW-BC [26]. We trained all methods for 200 epochs and evaluated the final checkpoints. As shown in Table II, we observe that SSDF+ACT achieves average success rates of 45% on these 5 tasks, which outperforms ACT by 6%. This increase in success rate shows that SSDF can select high-quality imperfect data, expand the original expert dataset, and thus improve the training of models.
@@ IV-C Ablation Study in Simulation
Impact of different self-supervised objectives. On the StackCube task, we evaluated SSDF with different combinations of the 3 self-supervised objectives in training step 1. In Table III, as we continue to increase the self-supervised objectives, SSDF can achieve higher success rates. This is because different self-supervised learning losses enable the model to learn versatile features, allowing it to compute quality scores for imperfect data accurately.
TABLE III: Ablation study on the different self-supervised objectives used in the training step 1 on the StackCube task.
StackCube 1 2 3 4 5 6 SSDF
MTP
TR
AA
Result 75.4 72.0 78.8 80.6 84.2 83.8 88.0
TABLE IV: Comparisons of different quality score threshold $\beta$ on the StackCube task.
@@
TABLE V: Comparisons of different similarity functions on the StackCube task.
@@
Impact of quality score threshold $\beta$. As shown in Table IV, on the StackCube task, we can observe that as the quality score threshold value $\beta$ increases, the reserved demonstrations become less and less. The highest success rate on the StackCube task is with $\beta$ as 0.90. When $\beta$ is less than 0.90, the performance increases as $\beta$ increases because more and more low-quality demonstrations are discarded. When $\beta$ is larger than 0.90, we argue that the performance drops as $\beta$ increases because more high-quality demonstrations are discarded, degrading the task performance.
Figure 4: Comparisons of different quantity and quality of the imperfect data. The x-axis represents the success rates of the behavior policies that collect the imperfect data. The y-axis shows the success rates of SSDF. The green and red lines represent the ratios of the expert and imperfect data, which are 1:1 and 1:3, respectively.
Impact of different similarity calculation functions. We examined $L_{2}$ distance and $Cosine$ similarity distance with the whole segment features denoted as “Seg.” and the last time step feature denoted as “Last” on the StackCube task. As shown in Table V, “Last+$L_{2}$” achieved the highest performance, while the results of using cosine similarity are relatively lower. We argue that the last time step of the trajectory segment is the most important step for predicting the next action, and the attention mechanism in the transformer has extracted the historical information.
Impact of the imperfect demonstration quantity and quality. For the StackCube task, we collected five different level imperfect demonstrations using different behavior policies with success rates of 0.00, 0.23, 0.44, 0.60, and 0.86, respectively . For each level, we collected imperfect demonstrations in two ratios, 1:1 and 1:3, where the former 1 represents the 300 expert demonstrations.
V Conclusion
This paper explores how to leverage imperfect demonstration to improve policy learning without reward information and online exploration. To this end, we propose a novel Self-Supervised Data Filtering framework (SSDF) that calculates accurate quality scores using a pre-trained transformer and then does weighted behavior cloning with the high-quality imperfect demonstrations. The extensive experimental results in simulation and real-world applications demonstrated that SSDF can accurately select high-quality imperfect demonstrations to boost the final performance.